LgswUnitDx work

LgswUnitDx installation

9:00: we go up to crane back the LGSW: ML leads the operation.

Landing_LGSW_on_cart.jpg Crane_strap_setup.jpg Flying_LGSW_over_primary.jpg Initial_removal_of_LGSW-DX.jpg

12:00: installation completed, we go back to control room and we switch on the LgswUnitDx.

New laser acquisition target

We measure the new FoV centers on Patrol cameras:

WFS X Y
Blue 179 146
Yellow 176 117
Red 177 127

  • Laser acquisition target tag: 20161015_192400

We also control and re-optimize the Patrol cameras clocking...

PockelsCellsTransmission test

TNs Plot
*DX: ['20161015_194932', ...] 20161015_194932_pcTransmission.jpg

Test with AdSec

We try to close the loop using the May reconstructor, LGSW only. We flatten the shell using flat 20160515_ARGOS_10modes

Best pupil position is with motors in:

WFS PM X PM Y
Blue 420 530
Yellow 300 370
Red 425 350

DX HEXA -5.16 -8.27 1.18 292 137
With IS we see the full pupils while with the cal source the bottom left part of yellow is vignetted (or fainter!) (left in figure below)

We compare the flux distribution with the one of May and we notice the same effect but on the Red at that time (right in figure below)

20161015_231901_FluxSubapMap.png 20160522_FluxSubapMap.png

LgswUnitDx work

LgswUnitSx Truth Sensing

With a mag=10 we test again the Truth Sensing closing the AO loop with the Pyramid and the LGSW. We move the Pyramid Board to null the differential focus between the Pyramid WFS and the LGSW.

TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20161015_173638 0.5,0.5,0.4 T 123.496 56.393 20161013_160000 -1 T F no dist
20161015_173704 0.5,0.5,0.4 T 104.349 48.4 20161013_160000 -1 T F
20161015_174323 0.6,0.5,0.4 T 51.8505 60.231 20161013_160000 -1 T F 1" dist
20161015_174341 0.6,0.5,0.4 T 63.5841 65.0478 20161013_160000 -1 T F
20161015_174400 0.6,0.5,0.4 T 52.9648 61.8985 20161013_160000 -1 T F
20161015_174500 0.6,0.5,0.4 T 65.9369 67.7154 20161013_160000 -1 T F
20161015_173638-20161015_174500_modalplot.png

APD calibrations

11:30 We connect the APD fibers to the AGW. Snapshot with APD in the dark: 20161015_193107 We start the dark calibration. We perfom a set of 4 seconds measures.
APD1 APD2 APD3 APD4
138 213 3106 275

Too much light on APD3, we go up and check the fibers.
APD1 APD2 APD3 APD4
140 300 400 300

It seems ok but then we try to measure the flat and we measured:
APD1 APD2 APD3 APD4
8000 12000 1000 11000

We change the APD3 fiber (FIBER2) with the SPARE2 and we measured a flat of:
APD1 APD2 APD3 APD4
10000 15000 9000 13000

We measured the dark and the flat with 40 seconds measures:
  APD1 APD2 APD3 APD4
dark [ph/s] 142 280 820 348
flat [%] 76 99 60 100
With the new dark and flat we close the AO loop with APD and LGSW. We use the Pyramid for the Truth Sensing with low flux (mag>18). We move the Pyramid Board to null the differential focus between the Pyramid WFS and the LGSW.

TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20161015_230338 0.5,0.5,0.4 T 2505.32 241 20161014_235200 -1 T F
20161015_231356 0.1,0.4,0.3 T 1202.54 33.9586 20161014_235200 -1 T F
20161015_232527 0.1,0.4,0.3 T 1388.8 34.0895 20161014_235200 -1 T F
20161015_232559 0.1,0.4,0.3 T 1308.09 34.0503 20161014_235200 -1 T F
20161015_234739 0.1,0.4,0.3 T 2191.83 42.8717 20161014_235200 -1 T F
20161015_234805 0.1,0.4,0.3 T 2192.53 43.2856 20161014_235200 -1 T F
20161015_235218 0.1,0.4,0.3 T 2179.27 43.2756 20161014_235200 -1 T F
20161015_235432 0.1,0.4,0.3 T 2692.42 47.34 20161014_235200 -1 T F
20161015_235736 0.1,0.4,0.3 T 1129.81 67.1931 20161014_235200 -1 T F
20161015_235757 0.1,0.4,0.3 T 1094.64 66.5177 20161014_235200 -1 T F
20161016_000127 0.1,0.4,0.3 T 1107.25 81.4256 20161014_235200 -1 T F
20161016_000148 0.1,0.4,0.3 T 1125.85 84.3245 20161014_235200 -1 T F
20161016_000204 0.1,0.4,0.3 T 1163.61 97.1494 20161014_235200 -1 T F
20161015_231356-20161016_000204_modalplot.png

The TS was able to converge to the same modal offset measured in the morning with a bright ngs.

20161015_modalOffset_PyrApd.png

ASM skip frame

After a few minutes of AO closed loop with the 1" disturbance we notice the Skip Frame flag and we check the temperature in the Housekeeper GUI of the ASM.

skipFrame.png
Topic revision: r4 - 28 Apr 2017, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback